Carter, Inc., a manufacturer of electrical supplies, has an ROE of 23.1 percent, a profit margin of 4.30 percent, and a total asset turnover ratio of 2.79 times. Its peer group also has an ROE of 23.1 percent but has out performed Carter with a profit margin of 5.16 percent and a total assets turnover ratio of 3 times. Calculate the Carter's equity multiplier and peer group equity multiplier. (Round answers to 2 decimal places, e.g.12.55.)
Answer #1

Equity multiplier \(=\frac{R O E}{P M \times T A T R}\)

PM : Profit Margin

TATR : Total assets turnover ratio

Carter's equity multiplier \(=\frac{23.1}{4.3 \times 2.79}=1.93\)

and peer group equity multiplier \(=\frac{23.1}{5.16 \times 3}=1.49\)